OpenStudy (brittany):

the graph of this quadratic function is a parabola. what is the equation of the axis of symmetry y=x^2+2x+1

OpenStudy (anonymous):

\[\frac{-b}{2a}\]

OpenStudy (anonymous):

hurray! \[-\frac{b}{2a}=-\frac{2}{2}=-1\] \[x=-1\] is your axis of symmetry